Thinfilms morning report 2/23/05

james haydon jhaydon at snf.stanford.edu
Wed Feb 23 15:31:47 PST 2005


           

Thinfilms Morning Report 2/23/2005

 

 

Gryphon:            ok

Innotec;            ion gauge controller not working I should have the 
parts tomorrow

STS Dep:            ok

Implanter:            noisy signal still but ion gauge problem is fixed 
if you look at the end station ion gauge I will try to get to the noise 
problem next week

Metalica:            problem with target tester will try to fix it next week

SCT:                ok

CPD:                ok                   

WbGen-CTB:            ok

WbGen-RHP:            ok

WbGen:-LHP:            ok

WbGaAs-RHP:            ok

WbGaSa-LHP:            ok

WbSilicide:            ok

WbDif: ok

WbNonmetal:            ok

WbNitride:            ok

WbMetal:            ok

Wbmiscres:            ok

WBOptres:            ok

WBSolvent:            ok

DI water:            ok

LN:                  ok

 

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/thinfilms/attachments/20050223/86b9910b/attachment.html>


More information about the thinfilms mailing list